# Hypergraphs as Weighted Directed Self-Looped Graphs: Spectral Properties, Clustering, Cheeger Inequality

### v1

@@ Abstract
Hypergraphs naturally arise when studying group relations and have been widely used in the field of machine learning. There has not been a unified formulation of hypergraphs, yet the recently proposed edge-dependent vertex weights (EDVW) modeling [7] is one of the most generalized modeling methods of hypergraphs, i.e., most existing hypergraphs can be formulated as EDVW hypergraphs without any information loss to the best of our knowledge. However, the relevant algorithmic developments on EDVW hypergraphs remain nascent: compared to spectral graph theories, the formulations are incomplete, the spectral clustering algorithms are not well-developed, and one result regarding hypergraph Cheeger Inequality is even incorrect. To this end, deriving a unified random walk-based formulation, we propose our definitions of hypergraph Rayleigh Quotient, NCut, boundary/cut, volume, and conductance, which are consistent with the corresponding definitions on graphs. Then, we prove that the normalized hypergraph Laplacian is associated with the NCut value, which inspires our HyperClus-G algorithm for spectral clustering on EDVW hypergraphs. Finally, we prove that HyperClus-G can always find an approximately linearly optimal partitioning in terms of Both NCut [^1] The NCut of the returned partition $\mathcal{N}$ and the optimal NCut of any partition $\mathcal{N}^{*}$ satisfy $\mathcal{N}\leq O(\mathcal{N}^{*})$. and conductance  [^2] The conductance of the returned partition $\Phi$ and the optimal conductance $\Phi^{*}$ satisfy $\Phi\leq O(\Phi^{*})$. Additionally, we provide extensive experiments to validate our theoretical findings from an empirical perspective.
@@ 1 Introduction
Higher-order relations are ubiquitous in nature, such as co-authorship [14, 48, 40], interactions between multiple proteins or chemicals [13, 47], items that are liked by the same person [49, 46], and interactions between multiple species in an ecosystem [20, 38]. Hypergraphs, extended from graphs, with the powerful capacity to model group interactions (i.e., higher-order relations), show extraordinary potential to be applied to many real-world tasks where the connections are beyond pair-wise. Therefore, hypergraphs have been used widely in recommendation systems [56, 36, 18], information retrieval [27, 53, 35] and link prediction [26, 12].
Hypergraphs modeled by edge-dependent vertex weights (EDVW) were necessitated in a recent work [7], with a motivating example that in citation networks, each scholar (i.e., vertex) may contribute differently to each co-authored publication (i.e., hyperedge). The authors show that hypergraphs with edge-independent vertex weights (EIVW) do not actually utilize the higher-order relations for the following two reasons. First, the hypergraph Laplacian matrix proposed by the seminal work [52], which serves as a basis of many follow-up algorithms, is equal to the Laplacian matrix of a closely related graph with only pair-wise relations. In this way, all the linear Laplacian operators utilize only pair-wise relationships between vertices [1]. Second, many hypergraph algorithms [37, 31, 4] are based on random walks [42, 9, 34], but it has been proved that for any EIVW hypergraph, there exists a weighted pair-wise graph on which a random walk is equivalent to that on the original hypergraph [7].
@@
edge/hyperedge weights $\surd$ $\surd$ $\surd$
vertex weights $\surd$ $\surd$ $\surd$
@@
Figure 1: Undirected graphs $\subset$ EIVW hypergraphs $\subset$ EDVW hypergraphs. Each undirected graph can be reformulated to EIVW hypergraph by regarding each pair-wise edge as a hyperedge; each EIVW hypergraph can be reformulated to EDVW hypergraph by setting each vertex’s weight to be the same across hyperedges, yet allowing different vertices to have different weights.
@@
In nature, ”EDVW hypergraph” is not a special case of hypergraphs, but a more generalized way to model hypergraphs (Figure 1). Any algorithm designed for EDVW hypergraphs, taking EDVW inputs, also works on typical (EIVW) hypergraphs by setting all the EDVW to 1 (i.e., independent w.r.t. edge). In other words, the properties and algorithms on EDVW-formulated hypergraphs can be applied to most hypergraphs.
@@
In this paper, we focus on further developing the incomplete yet fundamental spectral theories for EDVW hypergraphs, with a straightforward application on spectral clustering, a.k.a., k-way global partitioning, where typically $k=2$. To be specific, k-way global partitioning aims to partition an entire graph into $k$ clusters, where the vertices in one cluster are densely connected within this cluster while having sparser connections to vertices outside this cluster. On the one hand, although the spectral theories and spectral clustering on graphs have been well studied [10], converting the hypergraphs to graphs and applying those methods may ignore the higher-order relations and result in sub-optimal results [44]. On the other hand, despite the advantage of EDVW modeling in terms of utilizing high-order relations, directly developing a spectral clustering algorithm on EDVW hypergraphs is still an open question. To this end, for the first time, we propose a provably linearly optimal spectral clustering algorithm on EDVW hypergraphs, together with theoretical analysis concerning the Rayleigh Quotient, Normalized Cut (i.e., NCut), and conductance. In the context of EDVW hypergraphs, we bridge the eigensystem of Laplacian with the NCut value through our proposed Rayleigh Quotient. The proposed algorithm can also be applied to EIVW hypergraphs by setting all the vertex weights to 1, thus works generally for all hypergraphs.
@@ 1.1 Main Results
(algebraic connections between hypergraph NCut, Rayleigh Quotient and Laplacian) Given any hypergraph in the EDVW formatting $\mathcal{H}=(\mathcal{V},\mathcal{E},\omega,\gamma)$ with positive edge weights $\omega(\cdot)>0$ and non-negative edge-dependent vertex weights $\gamma_{e}(\cdot)$ for any $e\in\mathcal{E}$, define Normalized Cut $NCut(\cdot)$, Volume of a vertex set $vol(\cdot)$, Rayleigh Quotient $R(\cdot)$, Laplacian $L$, and stationary distribution matrix $\Pi$ as Definition 10, 8, 5, 13, and 4. For any vertex set $\mathcal{S}\subseteq\mathcal{V}$, we define a $|\mathcal{V}|$-dimensional vector $x$ such that,
$\begin{split}x(u)&=\sqrt{\frac{vol(\bar{\mathcal{S}})}{vol(\mathcal{S})}},\leavevmode\nobreak\ \forall\,\,u\in\mathcal{S},\\ x(\bar{u})&=-\sqrt{\frac{vol(\mathcal{S})}{vol(\bar{\mathcal{S}})}},\leavevmode\nobreak\ \forall\,\,\bar{u}\in\bar{\mathcal{S}}.\\ \end{split}$ (1)
@@
This is the first work regarding the Rayleigh Quotient on hypergraphs. Inspired by this Theorem, we develop a spectral clustering algorithm HyperClus-G to optimize the NCut value by loosing the combinatorial optimization constraint.
@@
(Hypergraph Spectral Clustering Algorithm) There exists a algorithm for hypergraph spectral clustering that can be applied to EDVW-formatted hypergraphs, and always returns approximately linearly optimal clustering in terms of Normalized Cut and conductance. In other words, approximately, the NCut of the returned partition $\mathcal{N}$ and the optimal NCut of any partition $\mathcal{N}^{*}$ satisfy $\mathcal{N}\leq O(\mathcal{N}^{*})$.
We name this algorithm as HyperClus-G. The pseudo code of HyperClus-G is given in Algorithm 1. Moreover, to extend the hypergraph spectral theory, for the first time we give a complete proof regarding the hypergraph Cheeger Inequality. In the mean time, by proving Theorem 3, the previous result on hypergraph Cheeger Inequality (Theorem 5.1 in [7]) is incorrect as it refers to the eigenvector of unnormalized hypergraph Laplacian.
@@
(Hypergraph Cheeger Inequality) Let $\mathcal{H}=(\mathcal{V},\mathcal{E},\omega,\gamma)$ be any hypergraph in the EDVW formatting with positive edge weights $\omega(\cdot)>0$ and non-negative edge-dependent vertex weights $\gamma_{e}(\cdot)$ for any $e\in\mathcal{E}$. Define $\Phi(\mathcal{H})=\mathop{\min}_{\mathcal{S}\subseteq\mathcal{V}}{\Phi(\mathcal{S})}$. Then the second smallest eigenvector $\lambda$ of the normalized hypergraph Laplacian $\Pi^{-\frac{1}{2}}L\Pi^{-\frac{1}{2}}$ satisfies
@@
In fact, this theorem shows that our HyperClus-G is also approximately linearly optimal in terms of conductance. In other words, the conductance of the returned cluster $\Phi$ and the optimal conductance $\Phi^{*}$ satisfy $\Phi\leq O(\Phi^{*})$. It is worth mentioning that the previous non-proved result in [7] regarding hypergraph Cheeger Inequality now can be proved by using normalized Laplacian instead of the conjecture of combinatorial Laplacian.
@@
Given the EDVW modeling, the relevant algorithmic development still remains in a nascent stage, which hinders the application of hypergraphs in many real-world scenarios. To this end, we first re-analyze the random walks on EDVW hypergraphs, then propose the HyperClus-G for hypergraph partitioning. Finally, we prove the approximation of normalized cut, as well as the upper bound of NCut and conductance.
The key insight from the previous work [7] is to model the hypergraphs similar to directed graphs through the equivalence of random walks. Unlike classical graph theory, such directed graphs are edge-weighted, node-weighted, and contain self-loops. In this work, inspired by the definitions of Rayleigh Quotient, NCut, boundary/cut, volume, and conductance in graphs, we develop these definitions in the context of EDVW hypergraphs. We show that Theorem 1 and Theorem 3, properties that hold for graphs, still hold for hypergraphs using our unified definitions. From Theorem 3, we can further prove that our proposed HyperClus-G is approximately linearly optimal in terms of both NCut and conductance.
@@
This paper is organized as follows. In Section 2, we introduce necessary notations and our definitions regarding hypergraphs. In Section 3, we introduce our definition of hypergraph Rayleigh Quotient and show its connection with the Laplacian and NCut. Then, we propose our HyperClus-G inspired from such connection. In section 4, we give complete proof regarding hypergraph Cheeger Inequality, then show the linear optimality of our HyperClus-G in terms of both NCut and conductance. In Section 5, we analyze the complexity of our algorithms. Finally, in Section 6, we prepare comprehensive experiments to validate our theoretical findings.
1.2 Other Related Works
Early Hypergraphs [6] typically model graph structures and do not allow node or hyperedge weights. Later formulations [52] start to allow hyperedge weights. Among various choices of modeling hypergraphs [6, 52, 15, 33], EDVW modeling [7] shows great generalization because it allows both hyperedge weights and node weights. Many hypergraph-related techniques have been proposed [17, 3, 30], while very few of them involve EDVW hypergraphs. Inhomogeneous hypergraph partitioning was proposed in [32]. Later on, authors of [33] proposed submodular hypergraphs, a special group of weighted hypergraphs, and analyzed their spectral clustering. A recent work [23] demonstrates how random walks with EDVW are used to construct the EDVW hypergraph Laplacian. However, it stops at the construction of the Laplacian and only uses partial information encoded in the Laplacian for clustering. Some recent works [2, 43, 11] study partitioning edge-colored hypergraphs. There are several research works [57, 54, 55] targeting EDVW hypergraph global partitioning, but they do not actually directly work on EDVW hypergraphs, but are based on submodular hypergraphs; They propose to construct a submodular hypergraph from the given EDVW hypergraph [33], then apply learning-based approaches to optimize the global partitioning objective. Several works also study specific applications of hypergraph clustering [28, 5].

### v2

@@ Abstract
Hypergraphs naturally arise when studying group relations and have been widely used in the field of machine learning. To the best of our knowledge, the recently proposed edge-dependent vertex weights (EDVW) modeling (Chitra & Raphael, 2019) is one of the most generalized modeling methods of hypergraphs, i.e., most existing hypergraph conceptual modeling methods can be generalized as EDVW hypergraphs without information loss. However, the relevant algorithmic developments on EDVW hypergraphs remain nascent: compared to the spectral theories for graphs, its formulations are incomplete, the spectral clustering algorithms are not well-developed, and the hypergraph Cheeger Inequality is not well-defined. To this end, deriving a unified random walk-based formulation, we propose our definitions of hypergraph Rayleigh Quotient, NCut, boundary/cut, volume, and conductance, which are consistent with the corresponding definitions on graphs. Then, we prove that the normalized hypergraph Laplacian is associated with the NCut value, which inspires our proposed HyperClus-G algorithm for spectral clustering on EDVW hypergraphs. Finally, we prove that HyperClus-G can always find an approximately linearly optimal partitioning in terms of both NCut [^1] The NCut of the returned partition $\mathcal{N}$ and the optimal NCut of any partition $\mathcal{N}^{*}$ satisfy $\mathcal{N}\leq O(\mathcal{N}^{*})$. and conductance [^2] The conductance of the returned partition $\Phi$ and the optimal conductance $\Phi^{*}$ satisfy $\Phi\leq O(\Phi^{*})$. Additionally, we provide extensive experiments to validate our theoretical findings from an empirical perspective. Code of HyperClus-G is available at https://github.com/iDEA-iSAIL-Lab-UIUC/HyperClus-G.
@@ 1 Introduction
Higher-order relations are ubiquitous, such as co-authorship (Feng et al., 2018; Yadati et al., 2019; Sun et al., 2021), interactions between multiple proteins or chemicals (Feng et al., 2021; Xia et al., 2022), items that are liked by the same person (Yang & Leskovec, 2015; Wu et al., 2021), and interactions between multiple species in an ecosystem (Grilli et al., 2017; Sanchez-Gorostiaga et al., 2019). Hypergraphs, extended from graphs, with the powerful capacity to model group interactions (i.e., higher-order relations), show extraordinary potential for real-world tasks beyond pair-wise connections. Therefore, hypergraphs have been used widely in recommendation systems (Zhu et al., 2016; Liu et al., 2022; Gatta et al., 2023), information retrieval (Huang et al., 2009; Zhu et al., 2015) and link prediction (Huang et al., 2020; Fan et al., 2022).
@@
edge and vertex weights $\surd$ $\surd$ $\surd$
@@
Table 2: A summary of our developed formulations of EDVW hypergraph spectral properties. Based on random walks on EDVW hypergraphs, we further develop the Rayleigh Quotient, NCut, boundary/cut, volume, and conductance, as well as their associations. "-" means not studied in Chitra & Raphael (2019).
Terminology/Property Graphs EDVW Hypergraphs
Chitra & Raphael (2019) Our work
Graph Definition $\mathcal{G}=(\mathcal{V},\mathcal{E})$ $\mathcal{H}=(\mathcal{V},\mathcal{E},\omega,\gamma)$. $\omega$ stores edge weight and $\gamma$ stores EDVW.
Random Walk Transition $P_{u,v}=\frac{\mathbbm{1}((u,v)\in\mathcal{E})}{degree(u)}$ $P_{u,v}=\sum_{e\in E(u)}\frac{\omega(e)}{d(u)}\frac{\gamma_{e}(v)}{\delta(e)}$, with row sum of 1.
Stationary Distribution $\phi P=\phi$ and $\sum_{u}\phi(u)=1$ $\phi P=\phi$ and $\sum_{u}\phi(u)=1$, diagonal matrix $\Pi(v,v)=\phi(v)$
Laplacian(s) L=DAL=D-A, $L_{sym}=D^{-\frac{1}{2}}LD^{-\frac{1}{2}}$ $L=\Pi-\frac{\Pi P+P^{T}\Pi}{2}$ $L=\Pi-\frac{\Pi P+P^{T}\Pi}{2}$, $L_{sym}=\Pi^{-\frac{1}{2}}L\Pi^{-\frac{1}{2}}$
Boundary Volume $|\partial S|=\sum_{u\in\mathcal{S},v\in\bar{\mathcal{S}}}\mathbbm{1}((u,v)\in\mathcal{E})$ - $|\partial S|=\sum_{u\in\mathcal{S},v\in\bar{\mathcal{S}}}\phi(u)P_{u,v}$
Vertex Set Volume $vol(S)=\sum_{u\in\mathcal{S}}degree(u)$ - $vol(S)=\sum_{u\in\mathcal{S}}\phi(u)$
Rayleigh Quotient $R(x)=\frac{\sum_{(u,v)\in\mathcal{E}}|x(u)-x(v)|^{2}}{\sum_{u\in\mathcal{V}}|x(u)|^{2}}$ - $R(x)=\frac{\sum_{u,v}|x(u)-x(v)|^{2}\phi(u)P_{u,v}}{\sum_{u}|x(u)|^{2}\phi(u)}$
Normalized Cut $(NCut)$ $(\frac{1}{vol(\mathcal{S})}+\frac{1}{vol(\bar{\mathcal{S}})})|\partial\mathcal{S}|$ - $NCut(\mathcal{S},\bar{\mathcal{S}})=(\frac{1}{vol(\mathcal{S})}+\frac{1}{vol(\bar{\mathcal{S}})})|\partial\mathcal{S}|$
Conductance $\Phi(\mathcal{S})=\frac{|\partial\mathcal{S}|}{\min(vol(\mathcal{S}),vol(\bar{\mathcal{S}}))}$ - $\Phi(\mathcal{S})=\frac{|\partial\mathcal{S}|}{\min(vol(\mathcal{S}),vol(\bar{\mathcal{S}}))}$
Algebraic Connections similar property as our theorem 1 - $NCut(\mathcal{S},\bar{\mathcal{S}})=\frac{1}{2}R(x)=\frac{x^{T}Lx}{x^{T}\Pi x}$, Theorem 1
Cheeger Inequality $\frac{\Phi(\mathcal{G})^{2}}{2}\leq\lambda_{L_{sym}}\leq 2\Phi(\mathcal{G})$ $\frac{\Phi(\mathcal{H})^{2}}{2}\leq\lambda_{L}\leq 2\Phi(\mathcal{H})$ $\times$ $\frac{\Phi(\mathcal{H})^{2}}{2}\leq\lambda_{L_{sym}}\leq 2\Phi(\mathcal{H})$, Theorem 3 $\surd$
Spectral Clustering well-developed - a new Algorithm 1 for EDVW hypergraphs
@@
Figure 1: Undirected graphs $\subset$ EIVW hypergraphs $\subset$ EDVW hypergraphs. Pair-wise edges are naturally hyperedges; each EIVW hypergraph can be reformulated to an EDVW hypergraph by setting each vertex’s weight to be the same across hyperedges, yet allowing different vertices to have different weights.
Hypergraphs modeled by edge-dependent vertex weights (EDVW) were necessitated in a recent work (Chitra & Raphael, 2019), with a motivating example that in citation networks, each scholar (i.e., node) may contribute differently to each co-authored publication (i.e., hyperedge). The authors show that hypergraphs with edge-independent vertex weights (EIVW) do not actually utilize the higher-order relations for the following two reasons. First, the hypergraph Laplacian matrix proposed by the seminal work (Zhou et al., 2006), which serves as a basis of many follow-up algorithms, is equal to the Laplacian matrix of a closely related graph with only pair-wise relations. In this way, all the linear Laplacian operators utilize only pair-wise relationships between vertices (Agarwal et al., 2006). Second, many hypergraph algorithms (Ma et al., 2018; Li et al., 2018; Carletti et al., 2020) are based on random walks (Tong et al., 2006), but it has been proved that for any EIVW hypergraph, there exists a weighted pair-wise graph on which a random walk is equivalent to that on the original hypergraph (Chitra & Raphael, 2019).
@@
In nature, EDVW hypergraphs are not a special case of hypergraphs, but a more generalized way to model hypergraphs by allowing flexible weight distribution, as shown in Figure 1. As typical (EIVW) hypergraphs can be reformulated to EDVW modeling, the properties and algorithms of EDVW-formulated hypergraphs can be applied to EIVW hypergraphs.
@@
In this paper, we focus on further developing the incomplete yet fundamental spectral theories for EDVW hypergraphs, with a straightforward application on spectral clustering, i.e., $k$-way global partitioning, where typically $k=2$. To be specific, $k$-way global partitioning aims to partition an entire graph into $k$ clusters, where the vertices in one cluster are densely connected within this cluster while having sparser connections to vertices outside this cluster. On the one hand, although the spectral theories and spectral clustering on graphs have been well studied, converting the hypergraphs to graphs (by connecting each node in a hyperedge) and applying those classic graph methods may ignore the higher-order relations and result in sub-optimal results (Wang & Kleinberg, 2024).
On the other hand, despite the advantage of EDVW modeling, directly developing a spectral clustering algorithm on EDVW hypergraphs is still an open question. To this end, for the first time, we propose a provably linearly optimal spectral clustering algorithm on EDVW hypergraphs, together with theoretical analysis concerning the Rayleigh Quotient, Normalized Cut (i.e., NCut), and conductance. In the context of EDVW hypergraphs, we bridge the eigensystem of Laplacian with the NCut value through our proposed Rayleigh Quotient. As discussed previously, the proposed algorithm can be applied to EIVW hypergraphs as well.
@@ 1.1 Main Results
(Algebraic connections among hypergraph NCut, Rayleigh Quotient, and Laplacian) Given any hypergraph $\mathcal{H}$ with vertex set $\mathcal{V}$ and hyperedge set $\mathcal{E}$ in the EDVW formatting, i.e., $\mathcal{H}=(\mathcal{V},\mathcal{E},\omega,\gamma)$ with positive edge weights $\omega(\cdot)>0$ and non-negative edge-dependent vertex weights $\gamma_{e}(\cdot)$ for any hyperedge $e\in\mathcal{E}$, we define Normalized Cut $NCut(\cdot)$, Volume of a vertex set $vol(\cdot)$, Rayleigh Quotient $R(\cdot)$, Laplacian $L$, and stationary distribution matrix $\Pi$ as Definition 16, 13, 13, 9, and 7. For any vertex set $\mathcal{S}\subseteq\mathcal{V}$, we define a $|\mathcal{V}|$-dimensional vector $x$ such that
@@
Then,
$NCut(\mathcal{S},\bar{\mathcal{S}})=\frac{1}{2}R(x)=\frac{x^{T}Lx}{x^{T}\Pi x}$ (2)
To the best of our knowledge, while prior works Jost & Mulas (2019); Mulas et al. (2022) studied the Rayleigh Quotient under the EIVW model, this is the first work that develops the Rayleigh Quotient for EDVW hypergraphs. Inspired by this Theorem, we develop a spectral clustering algorithm HyperClus-G to optimize the NCut value by relaxing the combinatorial optimization constraint.
@@
(Hypergraph Spectral Clustering Algorithm) There exists an algorithm for hypergraph spectral clustering that can be applied to EDVW-formatted hypergraphs, and always returns approximately linearly optimal clustering in terms of Normalized Cut and conductance. In other words, approximately, the NCut of the returned partition $\mathcal{N}$ and the optimal NCut of any partition $\mathcal{N}^{*}$ satisfy $\mathcal{N}\leq O(\mathcal{N}^{*})$.
We name this algorithm as HyperClus-G, whose pseudo code is given in Algorithm 1 with complexity analyzed in Appendix B. Moreover, to extend the hypergraph spectral theory, for the first time, we give complete proof regarding the hypergraph Cheeger Inequality. In the meantime, by proving Theorem 3, the previous result on hypergraph Cheeger Inequality (Theorem 5.1 in (Chitra & Raphael, 2019)) is fixed and upgraded.
@@
(Hypergraph Cheeger Inequality) Let $\mathcal{H}=(\mathcal{V},\mathcal{E},\omega,\gamma)$ be a hypergraph in the EDVW formatting. Define $\Phi(\mathcal{H})=\mathop{\min}_{\mathcal{S}\subseteq\mathcal{V}}{\Phi(\mathcal{S})}$. Then the second smallest eigenvector $\lambda$ of the normalized hypergraph Laplacian $\Pi^{-\frac{1}{2}}L\Pi^{-\frac{1}{2}}$ satisfies
@@
In fact, this theorem shows that our HyperClus-G is also approximately linearly optimal in terms of conductance. In other words, the conductance of the returned cluster $\Phi$ and the optimal conductance $\Phi^{*}$ satisfy $\Phi\leq O(\Phi^{*})$. As shown in Table 2, we bridge the gap between the previous work Chitra & Raphael (2019) to broader spectral applications by further developing spectral theories on EDVW formulation.
@@
The technical map is summarized in Figure 2. The key insight from the previous work (Chitra & Raphael, 2019) is to model the hypergraphs similar to directed graphs through the equivalence of random walks. Unlike classical graph theory, such directed graphs are edge-weighted, node-weighted, and contain self-loops. In this work, inspired by the definitions of Rayleigh Quotient, NCut, boundary/cut, volume, and conductance in graphs, we develop these definitions in the context of EDVW hypergraphs. We show that Theorem 1 and Theorem 3, properties that hold for graphs, still hold for hypergraphs using our unified definitions. From Theorem 3, we can further prove that our proposed HyperClus-G is approximately linearly optimal in terms of both NCut and conductance.
@@
In Section 2, we introduce notations regarding EDVW hypergraphs. In Section 3, we introduce our definition of hypergraph Rayleigh Quotient and show its connection with the Laplacian and NCut. Then, we propose our HyperClus-G inspired by such a connection. In section 4, we prove the Cheeger Inequality of EDVW hypergraph, then show the optimality of our HyperClus-G. Finally, Section 5 presents comprehensive experiments to validate our theoretical findings.
{forest}
Figure 2: Logical flow of our work. "A->B" means A is required for developing B.